The momentum p of a particle changes with time t according to the relation dp/dt=(10 N) + (2 N/s)t. If the momentum is zero at t = 0, what will the momentum t be at t = 10 s ?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. . dp/dt (10 N) + (2 N/S)t Momentum is zero at t = 0 ∴ momentum at t = 10 sec will be Dp = [(10 N) + 2Ns t] dt ∫_0^p▒dp = ∫_0^10▒〖10 dt〗 + ∫_0^10▒〖(2tdt)〗 = 10〖t]〗_0^10 + ├ 2 t^2/2]_0^10 = 200 kg m/s.